The UFS estimates that they will need $450000 in one and a half year's time and another $720000 in two and a half years' time to cover the expected cost of providing each registered student with a Kovsie dairy. If interest is calculated at 6.4% per annum, compounded semi-annually, how much must the UFS invest today in order to cover the expected future costs?
1
Expert's answer
2020-04-21T12:24:29-0400

"\frac{450 000}{(1+\frac{6.4}{2})^{1.5\times2}}+\frac{720 000}{(1+\frac{6.4}{2})^{2.5\times2}}=409 424.117+615083.403=1 024 507.52"